OpenStudy (anonymous):

A gem store sells beads made of amber and quartz. For 4 amber beads and 4 quartz beads, the cost is $46.00. For 1 amber bead and 3 quartz beads, the cost is $14.50. Find the price of each type of bead.

OpenStudy (anonymous):

Pricw of amber bead x price of quartz bead y 4x+4y=46 x+3y=14.5

OpenStudy (anonymous):

solve for x and y

OpenStudy (anonymous):

Problem: A gem store sells beads made of amber and quartz. For 4 amber beads and 4 quartz beads, the cost is $46.00. For 1 amber bead and 3 quartz beads, the cost is $14.50. Find the price of each type of bead. This is a typical simultaneous equation problem. Let's let "A" represent the cost of one amber bead and "Q" represent the cost of one quartz bead. So we have two equations: 4A + 4Q = 46 A + 3Q = $14.50 I assume you know how to solve simultaneous equations? If not, here's one thing you can do: multiply the second equation (both sides) by -4; then add that equation to the first to eliminate the "A" variable; then solve for "Q". Then plug that Q value into the second equation and solve for A. You'll then have the price for each bead (Q and A). Good luck!